1. Write an equation of the line that passes through the given two points ( -3 , 5 ), ( 41 , -127 )
2. Write the slope of the line passing through the two points ( -4 , -1 ), ( -6 , 6
3.Write the slope of the line passing through the two points ( 7 , -3 ), ( -3 , 5 )
4. Solve each system of equations using the elimination method5x + 6y = -75x - 3y = 146
5.Write the slope-intercept form of the equation 3y = -15 + 3x

Answer by Alan3354(69443)

Write an equation of the line that passes through the points (-3,5), (41,127)
------------
Given 2 points example.
Find the equation of the line thru the points (2,1) and (3,5)
---------------------
This is a 2 step process. First find the slope of the line thru the points.
slope, m = diffy/diffx
m = (5-1)/(3-2)
m = 4
---------
Now use y = mx + b with either point to find b, the y-intercept.
y = mx + b
5 = 4*3 + b
b = -7
-------
y = mx + b
y = 4x - 7 is the answer.
------------------------
Another approach using determinants:
|x y 1|
|2 1 1| = 0
|3 5 1|
--------
x(1-5) - y(2-3) + 1(10-3) = 0
-4x + y + 7 = 0
y = 4x - 7
---------
For further assistance, or to check your work, email me via the thank you note
or at Moral Loophole@aol.com